40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1900     Entry time: Fri Aug 14 02:57:46 2009
Author: Clara 
Type: Update 
Category: PEM 
Subject: Redo of the Huddle Test 

I put all three seismometers and all six accelerometers together in the foam box with peanuts. Three of the accelerometers are facing in the x-direction and three are in the y-direction. Both Guralps are aligned on the NS axis and the Ranger is pointing vertically.

**EDIT: The accelerometers are in the x and z directions, not x and y. Sorry, I was sleepy when I wrote this.**

One of the accelerometers was refusing to show anything, and after a few hours of checking connections and swapping cables, I discovered that someone had unplugged the cable from the ADC. A quick glance in the dataviewer shows that the channel has been unplugged since about 3 in the afternoon on August 8th (Saturday). So... obviously all the accelerometer measurements made with that channel since then did not actually get recorded. Yay.

Anyway, as of 2:45, everything is working and taking data. Clearly we're not getting a full night's worth... hopefully that's okay.

ELOG V3.1.3-